OpenStudy (anonymous):

Integral calc. is not my friend. Wondering if there is an easier way to solve the following integral than knowing/deriving the antiderivatives of csc(x) and sec(x)...

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ sinx+cosx }{ \sin(2x)}\]

OpenStudy (anonymous):

\[\int\limits_{}{}\frac{ sinx+cosx }{ 2sinxcosx }\] \[\int\limits_{}^{} \frac{ sinx }{ 2sinxcosx }+\int\limits_{}^{}\frac{ cosx }{ 2sinxcosx }\] \[1/2\int\limits_{}^{}(cscx+secx)\]

OpenStudy (anonymous):

dx everywhere

OpenStudy (anonymous):

write sin2x as 1- (sinx - cos x)^2 and substitute sin x- cos x as t and sin x +cos x will become dt